OpenStudy (anonymous):

What is a simpler form of the radical expression? 3^sq root of 125x^21y^24 5x^7|y^8| 15x^21|y^24| 15|x^21|y^24 5x^7y^8

OpenStudy (anonymous):

Do you mean \[3^{\sqrt{125x^{21}y^{24}}}\]

OpenStudy (anonymous):

or this \[\sqrt[3]{125x^{21}y^{24}}\] ?

OpenStudy (anonymous):

The second one

OpenStudy (anonymous):

does splitting it up like this: \[\sqrt[3]{125}\sqrt[3]{x^{21}}\sqrt[3]{y^{24}}\] help you out? Also, \[\sqrt[b]{x^{a}} = x^{\frac{a}{b}}\]

OpenStudy (anonymous):

Im a little lost.

OpenStudy (n00bstyle):

(a*b*c)^2 = a^2 * b^2 * c^2

OpenStudy (anonymous):

^ Is that the formula?

OpenStudy (anonymous):

yes, you can use any number, not just 2

OpenStudy (anonymous):

which is why I suggested \[\sqrt[b]{x^{a}}=x^{\frac{a}{b}}\]

OpenStudy (anonymous):

so the whole thing becomes \[(125x^{21}y^{24})^{\frac{1}{3}}\]

OpenStudy (anonymous):

and then you can use N00bstyle's "formula" to solve

OpenStudy (n00bstyle):

\[\sqrt[3]{125}=125^{1/3}=(5^{3})^{1/3}=5\]

OpenStudy (n00bstyle):

This is one example, you can apply this to the other two...let me know if you manage to do that

OpenStudy (anonymous):

so 125x^21=1953125?

OpenStudy (anonymous):

@_@ im too confused

OpenStudy (n00bstyle):

Okay first rule when you have multiplications under a square root: \[\sqrt{ab}=\sqrt{a}\sqrt{b}\]

OpenStudy (n00bstyle):

Do you understand this?

OpenStudy (anonymous):

uhmm, yes

OpenStudy (n00bstyle):

Okay, you have a multiplication of 125*x^21*y^24

OpenStudy (anonymous):

oh, okay! Im getting it.

OpenStudy (anonymous):

But how would multiplying make a simpler form?

OpenStudy (n00bstyle):

\[\sqrt[3]{125}*\sqrt[3]{x^{21}}*\sqrt[3]{y^{24}}\]

OpenStudy (n00bstyle):

NOW you apply the rule which aachem said: \[\sqrt[y]{x}=x^{1/y}\]

OpenStudy (anonymous):

?

OpenStudy (n00bstyle):

To be more precise: \[\sqrt[y]{a^{x}}=a^{x/y}\]

OpenStudy (n00bstyle):

So when you rewrite \[\sqrt[3]{125}\] You get: \[125^{1/3}\]

OpenStudy (n00bstyle):

If you don't understand, just tell me :)

OpenStudy (anonymous):

so 125*.5=62.5

OpenStudy (n00bstyle):

that's correct, but that's not what you want to know :)

OpenStudy (anonymous):

what do i need to know?

OpenStudy (anonymous):

D:

OpenStudy (n00bstyle):

Read the first message aachem wrote you

OpenStudy (n00bstyle):

Sorry, third message

OpenStudy (anonymous):

OH! THE SQUARE ROOT!

OpenStudy (n00bstyle):

He splits up all the multiplications

OpenStudy (n00bstyle):

afterwards, you just need to apply the rule he gave you. Later in the text, I gave you an example about the cubic root of 125

OpenStudy (anonymous):

and multiply's each number by the power of 3?

OpenStudy (n00bstyle):

he takes the cubic root of every numbers yes

OpenStudy (n00bstyle):

that's what the rule says: \[\sqrt{ab}=\sqrt{a}\sqrt{b}\]

OpenStudy (n00bstyle):

This also applies to: \[\sqrt[3]{abc}=a^{1/3}b^{1/3}c^{1/3}\]

OpenStudy (anonymous):

...

OpenStudy (n00bstyle):

Let me write it for you, it will give you the answer, but it's important that you know what happens okay

OpenStudy (anonymous):

Yes

OpenStudy (n00bstyle):

To me it's important you know how the rules can be applied. When you know the rules you can apply it to every equation...

OpenStudy (anonymous):

k

OpenStudy (n00bstyle):

\[\sqrt[3]{125x^{21}y^{24}}=\sqrt[3]{125^{1} x^{21}y^{24}}=\sqrt[3]{125}\sqrt[3]{x^{21}}\sqrt[3]{y^{24}}\] This is the first part of the reformulation of the equation

OpenStudy (anonymous):

So you split them

OpenStudy (anonymous):

Then divide?

OpenStudy (n00bstyle):

\[125^{1/3}(x^{21})^{1/3}(y^{24})^{1/3}\]

OpenStudy (anonymous):

ok, you divide by 3

OpenStudy (anonymous):

OH!

OpenStudy (n00bstyle):

Now I applied the 'second' rule: \[\sqrt[y]{a^{x}}=a^{x/y}\]

OpenStudy (n00bstyle):

that's why there are all 1/3 in the equation

OpenStudy (anonymous):

Its making sense now!

OpenStudy (n00bstyle):

Now you will have to apply rules of multiplying exponents...do you know that?

OpenStudy (zehanz):

OK, I will give it a try! You must remember that to be able to easily calculate a third root, you'll need a number that has 3 same factors. E.g.:\[\sqrt[3]{125}=5\]because \[125=5^3\]You have to simplify:\[\sqrt[3]{125x^{21}y^{24}}=\sqrt[3]{5^3x^{21}y^{24}}\]Now, part of your problem is already done, the thrird root of 125 is 5, so we have left:\[5\sqrt[3]{x^{21}y^{24}}\]Oh, but now you can see that all these factors x can also be grouped in three same powers:\[x^{21}=x^7x^7x^7\]and \[y^{24}=y^8y^8y^8\]So we can calculate the third root of these numbers, just like we did for 125:\[\sqrt[3]{x^7x^7x^7}=x^7\]and\[\sqrt[3]{y^8y^8y^8}=y^8\]So the end result must be:\[\sqrt[3]{125x^{21}y^{24}}=5x^7y^8\] The only thing you must remember is: make 3 equal groups of powers. What if you had to do this:\[\sqrt[3]{x^{22}}\]Obviously, 22 is not a multiple of 3, but we will do as much as we can: \[\sqrt[3]{x^{22}}=\sqrt[3]{x^7x^7x^7x}=x^7\sqrt[3]{x}\]Everything goes as expected, only the last x is stuck under the radical. Bad luck!

OpenStudy (n00bstyle):

so: \[(x^{21})^{1/3}=x^{7}\]

OpenStudy (anonymous):

and 1/3*125=41.666666666666666666666666666667 41.666666666666666666666666666667 squared is 6.4?

OpenStudy (anonymous):

so if its 6.4 with x^7 it must be 5x^7y^8?

OpenStudy (anonymous):

0_0

OpenStudy (n00bstyle):

I don't know what you mean by 41.6666666666 squared is 6.4. When you square something, you have 41.66666^2 = 41.66666 * 41.66666 = a big number

OpenStudy (anonymous):

oops, i mean square root

OpenStudy (n00bstyle):

ah okay, I see too

OpenStudy (n00bstyle):

5x^7y^8 is idd the answer, but you will have to take a look at: exponents rules and mutiplication with square roots

OpenStudy (anonymous):

Defiantly

OpenStudy (anonymous):

whew

OpenStudy (n00bstyle):

Just try to remember: \[\sqrt[y]{a^{x}} = a^{x/y}\] So for example: \[\sqrt[3]{8}=8^{1/3}=(2^{3})^{1/3}=2^{3*1/3}=2^{1}=2\]

OpenStudy (anonymous):

Yes, il take notes just in case I forget, Thank you!

OpenStudy (n00bstyle):

You'll get used to in time...just practise a lot ;)

OpenStudy (anonymous):

Thanks for your patience! I can be allot to deal with!

OpenStudy (n00bstyle):

Ha, you're welcome, good luck.